OpenStudy (anonymous):

im having trouble with this one. Find the 15th term of the sequence in which a1 = -11 and the common difference is -6. with the equation i have im getting -795, and thats not right

OpenStudy (ash2326):

nth term of an arithmetic sequence is given as \[a_n=a_1+(n-1)d\] \(a_1\)= first term d= common difference here \(a_1\)=-11 d=-6 n=15 \[a_{15}=a_1+(15-1)d\] \[a_{15}=-11+(14)\times (-6)\] Now you evaluate and check if it matches the answer?

OpenStudy (anonymous):

the answer would be -95 then

OpenStudy (ash2326):

Yeah:D

OpenStudy (anonymous):

awesome! thanks a lot!!! now i was given an equation earlier that had me divide by 2 in the begining, does that mean i have the incorrect answers?

OpenStudy (ash2326):

What equation you had? Can you post the equation?

OpenStudy (anonymous):

s=n/2(2a+(n-1)d)

OpenStudy (ash2326):

That's for the sum of the sequence. It's correct but not applicable here

OpenStudy (anonymous):

ok i see, thanks again ! :D

OpenStudy (ash2326):

Welcome:D